Grade 11Physical Chemistry

1. A 6.85 grams sample of the hydride Sr(OH)2.xH2O is dried in a oven to give 3.13 g of anhydrous 2 Sr(OH) .
What is the value of x? (At. wt. Sr = 87.6)

divya
10 Years agoGrade 11

1 Answer

Ashutosh Mohan Sharma
10 Years ago
Reduction in weight = 6.85 – 3.13 = 3.72 g
the only reason for reduction in weight is evaporation of water.
therefore amount of water evaporated = 3.72 g
therefore moles of water = 3.72/ 18 = 2.06 moles
therefore value of x is aproximately 2.